# The sphere complex of a locally finite graph

### v1

@@ 1 Introduction
Essentially all of the work so far in studying mapping class groups of infinite type graphs has been in understanding the basic algebraic and topological properties of these groups, as well as their coarse geometry [algom-kfir_groups_2021] [domat_coarse_2023] [domat_generating_2023]. Most of it is focused on studying the pure mapping class group PMap$(\Gamma)$, the subgroup which fixes the ends of the $\Gamma$.
In this paper, we introduce a geometrically defined complex that Map$(\Gamma)$ acts on naturally. Namely, given a locally finite graph $\Gamma$, we consider the associated doubled handlebody with punctures $M_{\Gamma}$. This is a $3$-manifold obtained by gluing two copies of a regular neighborhood of $\Gamma$ in ${\mathbb{R}}^{3}$ along their common surface boundary by the identity map (see Definition 2.3 for more details). Then the sphere complex $\mathcal{S}(M_{\Gamma})$ is the simplicial complex whose $k$-simplices are given by isotopy classes of $k+1$ essential non-peripheral smoothly embedded spheres in $M_{\Gamma}$ which can be realized disjointly. The mapping class group Map$(M_{\Gamma})$ of $M_{\Gamma}$, which is the group of orientation preserving diffeomorphisms of $M_{\Gamma}$ up to isotopy, acts naturally on this complex.
@@
Let $\Gamma$ be a locally finite graph, and let $M_{\Gamma}$ denote its associated doubled handlebody with punctures.
@@
    where $\mathrm{Twists}(M_{\Gamma})$ denotes the subgroup of $\mathrm{Map}(M_{\Gamma})$ generated by sphere twists on sphere systems of $M_{\Gamma}$. Further, as long as $\Gamma$ is not a rank $1$ graph with 00 or $1$ end(s), $\mathrm{Twists}(M_{\Gamma})$ is precisely the kernel of the action of $\mathrm{Map}(M_{\Gamma})$ on $\mathcal{S}(M_{\Gamma})$, so there is an induced faithful action of $\mathrm{Map}(\Gamma)$ on $\mathcal{S}(M_{\Gamma})$.
@@
Here, a sphere system is a smooth proper embedding of (possibly infinitely many) disjoint essential non-peripheral $2$-spheres in $M_{\Gamma}$. Unless otherwise stated, we assume that no pair of spheres are isotopic. A sphere in a $3$-manifold $M$ is essential if it does not bound a $3$-ball, and non-peripheral if it does not bound a once punctured $3$-ball and it is not isotopic to a boundary component. A sphere system is simple if its complement is simply connected (this is slightly more restrictive than the normal definition of simple, where the complement can be disconnected, but each component is simply connected).
The assumption in Theorem 1.1(1) that $\Gamma$ is not a rank $1$ graph with 00 or $1$ end(s) is because in these cases, $\mathcal{S}(M_{\Gamma})$ is a single vertex, and $\text{Map}(\Gamma)\cong{\mathbb{Z}}/2$, so the induced action is not faithful.
@@
The group $\mathrm{Twists}(M_{\Gamma})$ is topologically isomorphic to $\Pi_{i=1}^{\text{rk}(\Gamma)}{\mathbb{Z}}/2$, with generating set given by sphere twists on a simple sphere system (allowing for infinite products). In particular, $\mathrm{Twists}(M_{\Gamma})$ is compact.
@@
Finally, in Subsection 8.2 we consider a particular collection of graphs known as translatable graphs, analgous to the translatable surface of Schaffer-Cohen [schaffer-cohen_graphs_2024]. These can be thought of as graphs which can be formed by taking a collection of copies of a fixed graph arranged like ${\mathbb{Z}}$ with neighboring copies wedge summed along a point. See Definition 8.12 for a more precise definition. For a particular subclass of these graphs (those with tame end space, see Definition 8.16), we have the following.
@@
Acknowledgements: The author thanks his advisor Christopher J. Leininger for his support and guidance throughout this project. He’d also like to thank George Domat, Hannah Hoganson, and Mladen Bestvina for discussions about their work and for their support. Thanks to Thomas Hill, Michael Kopreski, Rebecca Rechkin, and George Shaji for helpful conversations on the topic while working on a project in parallel to this one. Finally, thanks to Sanghoon Kwak for helpful comments and corrections on an earlier version of the paper.
@@ 2.2 Graphs and their doubled handlebodies
Let $\Gamma$ be a locally finite connected graph (every graph considered in this paper will be locally finite and connected, unless otherwise specified). Typically, $\Gamma$ will be noncompact. We can consider the space of ends of $\Gamma$. The space of ends accumulated by loops of $\Gamma$, denoted $E_{\ell}(\Gamma)$, is the closed subset of $E(\Gamma)$ consisting of those ends whose neighborhoods in $\Gamma$ are all of infinite rank. Note that the elements of $E_{\ell}$ are called non-$\infty$-stable ends in [ayala_proper_1990] and ends accumulated by genus in [algom-kfir_groups_2021].
@@
The doubled handlebody $M_{\Gamma}$ of $\Gamma$ is the union of all the pieces of the vertices of $\Gamma$, with two boundary spheres quotiented together if the corresponding vertices are connected by an edge in $\Gamma$.
@@
Fix a maximal sphere system $\mathcal{P}$ in $M_{\Gamma}$ (i.e. a sphere system so every other sphere either intersects or is isotopic to some component of it). If $\Gamma$ is of infinite type, then $\mathcal{P}$ is infinite. Such a maximal system splits $M_{\Gamma}$ into a collection $\{P_{k}\}$ of copies of $M_{0,3}$ (sometimes one has to replace a boundary component of a copy of $M_{0,3}$ with a puncture corresponding to an end of $M_{\Gamma}$).
@@ 2.3 The Sphere Complex
An isotopy of $M_{\Gamma}$ preserves the ends of $M_{\Gamma}$, so if two spheres are isotopic they must induce the same partition of $E(\Gamma)$. On the other hand, suppose $S$ and $S^{\prime}$ induce the same partition of $E(\Gamma)$. As $S$ and $S^{\prime}$ are compact, we can choose a finite number of disjoint spheres $T_{1},\ldots,T_{n}$ which are disjoint from $S$ and $S^{\prime}$ so that $S,S^{\prime}$ lie inside the submanifold bounded by $T_{1},\ldots,T_{n}$, and so that this submanifold is compact. As $S$ and $S^{\prime}$ induce the same partition of $E(\Gamma)$, they induce the same partition of the spheres $T_{1},\ldots,T_{n}$. But isotopy classes of spheres in manifolds homeomorphic to $M_{0,n}$ are determined by their partitions of boundary spheres, see Lemma 9 of [bering_iv_finite_2024]. ∎
Fix a sphere system $Q$ and an embedded sphere $S$ in normal form with respect to $Q$. We may identify an innermost disc $D$ of this intersection, which is a disc in a component of $Q$ so that $D$ does not contain any other component of $D\cap Q$. The surgery of $S$ along $D$ is the pair of spheres which result from cutting $S$ along $D$, and then attaching copies of $D$ to each of the two components. The resulting spheres are embedded and essential, though not always non-peripheral.
@@
Fix a basepoint $p\in\partial M_{n,1}$. We need one more auxiliary complex, which we will denote by $\widehat{\mathcal{S}}_{ns}(M_{n,1})$. It is the flag complex whose vertex set agrees with $\mathcal{S}_{ns}(M_{n,1})$. If $n\geq 3$, then there is an edge between two vertices $S$ and $S^{\prime}$ if $\pi_{1}(M_{n,1}\setminus(S\cup S^{\prime}),p)$ is nontrivial. If $n=2$, then there is an edge between vertices $S$ and $S^{\prime}$ if $\pi_{1}(M_{n,1}\setminus S,p)$ and $\pi_{1}(M_{n,1}\setminus S^{\prime},p)$ generate $\pi_{1}(M_{n,1},p)$. Note that for any $n\geq 2$, an edge in $\mathcal{S}_{ns}(M_{n,1})$ between vertices $S$ and $S^{\prime}$ is not in $\widehat{\mathcal{S}}_{ns}(M_{n,1})$ if and only if $S$, $S^{\prime}$, and $\partial M_{n,1}$ bound a copy of $M_{0,3}$.
@@
    Fix two disjoint vertices $S$ and $S^{\prime}$. If $S$ and $S^{\prime}$ are not joined by an edge in $\widehat{\mathcal{S}}_{ns}(M_{n,1})$, then it follows from the note above that the complementary component of $S\cup S^{\prime}$ not containing $p$ is homeomorphic to $M_{n-1,2}$. As $n\geq 2$, there is a nonseparating sphere $S^{\prime\prime}$ in this component.
@@ 3.1 Mapping class groups of doubled handlebodies and graphs
where $\ell(t)\cdot x$ denotes the action of $\ell(t)\in\text{SO}(3)$ on $x\in S\cong S^{2}\subset{\mathbb{R}}^{3}$. We denote by $\text{Twists}(M_{\Gamma})$ the subgroup of $\text{Map}(M_{\Gamma})$ generated by the sphere twists of $M_{\Gamma}$ (allowing for infinite products).
@@ 3.2 Constructing the homomorphism
We now construct the homomorphism from $\text{Map}(M_{\Gamma})$ to $\text{Map}(\Gamma)$ which appears in Theorem 1.1
@@ 4 Short exact sequence
In this section, we give a proof of the first part of Theorem 1.1. The proof of the second part requires Theorem 1.2, and will be done in Section 5. We split the proof into two cases. The first case, where we assume $\Gamma$ has rank 00 or rank at least $2$, will be handled in Lemmas 4.2 and 4.3. The rank $1$ case will be handled in Lemma 4.4.
@@
We will apply this, along with the following lemma, to obtain Theorem 1.1 in the case when $\Gamma$ has rank 00 or rank at least $2$. This lemma is due to Laudenbach in the case that $\Gamma$ is compact [laudenbach_topologie_1974]. Our proof is based on the proof of the same result by Brendle–Broaddus–Putman. [brendle_mapping_2023].
@@
We lift to the universal cover $\widetilde{\Gamma}\subset\widetilde{M_{\Gamma}}$ with covering map $\pi:\widetilde{M_{\Gamma}}\to M_{\Gamma}$ and with a fixed lift $\widetilde{x_{0}}$.. Let $\tilde{f}:\widetilde{M_{\Gamma}}\to\widetilde{M_{\Gamma}}$ be the unique lift of $f$ fixing $\widetilde{x_{0}}$. We also denote by $\widetilde{i}:\widetilde{\Gamma}\to\widetilde{M_{\Gamma}}$ and $\widetilde{r}:\widetilde{M_{\Gamma}}\to\widetilde{\Gamma}$ the inclusion and retraction maps so that $\pi\circ\widetilde{i}=i\circ\pi$ and $\pi\circ\widetilde{r}=r\circ\pi$. To show that $f$ acts trivially on $\pi_{2}(M_{\Gamma},x_{0})$, it suffices to show that $\widetilde{f}$ acts trivially on $\pi_{2}(\widetilde{M_{\Gamma}},\widetilde{x_{0}})\cong H_{2}(\widetilde{M_{\Gamma}})$, by the Hurewicz Theorem.
@@
The next lemma gives a more explicit form of elements of $\text{ker}(\Psi)$. First, we make the following definition. For a fixed choice of $n,s\geq 0$, we make a choice of core spheres for $M_{n,s}$. That is, we consider $M_{n,s}$ as a connect sum of $n$ copies of $S^{2}\times S^{1}$’s, with $s$ balls removed. In each $S^{2}\times S^{1}$ factor, there is a unique nontrivial embedded copy of $S^{2}$. The collection of these spheres are the core spheres of $M_{n,s}$. For any piece of $M_{\Gamma}$ homeomorphic to $M_{n,s}$, we identify the core spheres of it via the identification of this piece with $M_{n,s}$. The collection of core spheres of $M_{\Gamma}$ is then the union of the core spheres of its pieces.
@@
Suppose $\Gamma$ has rank 00 or rank at least $2$. Fix $f\in\text{Diff}^{+}(M_{\Gamma})$ acting trivially on $\pi_{2}(M,x_{0})$ for some $x_{0}\in M$. Then $f$ is isotopic to a product of sphere twists on the core spheres and the boundary components of the pieces of $M_{\Gamma}$. In particular, $\mathrm{ker}(\Psi)$ is a compact abelian subgroup of $\mathrm{Map}(M_{\Gamma})$.
@@
Let $\mathcal{S}=\sqcup_{i}(\partial K_{i}\cup\mathcal{C}_{i})$. By Theorem 4.1, $f$ acts trivially on each isotopy class of each component of $\mathcal{S}$, The goal is to produce an ambient isotopy of $M_{\Gamma}$ so that after applying this isotopy to $f$, every component of $\mathcal{S}$ is actually fixed by $f$. Note that we cannot apply Theorem 4.1 directly, as $\mathcal{S}$ has infinitely many elements. However, we will see that it can be inductively applied with care to achieve the desired result.
@@
Let $\widetilde{M_{\Gamma}}$ denote the universal cover of $M_{\Gamma}$ with $\pi:\widetilde{M_{\Gamma}}\to M_{\Gamma}$ the covering map. We can lift the homotopy between $S$ and $f_{n}(S)$ to one between lifts of these two spheres in $\widetilde{M_{\Gamma}}$. Let us denote these lifts by $T_{1}$ and $T_{2}$. By Theorem 4.1, they are isotopic as well. Thus these spheres lie in the same component $U$ of $\widetilde{M_{\Gamma}}\setminus\pi^{-1}(K_{n})$, or else there is no way they could induce the same partition of $E(\widetilde{M_{\Gamma}})$ (using the assumption that $K_{i}$ is not homeomorphic to $M_{0,2}$), contradicting Lemma 2.13. But then they also induce the same partition of $E(U)$, thinking of its boundary as an end. Again by Lemma 2.13, they are isotopic in $U$. This restricts to a homotopy from $T_{1}$ to $T_{2}$ whose image lies in $U$.
@@
Then it follows by Lemma 4.2 that $\text{Ker}(\Psi)$ is generated by compositions of sphere twists on a proper sphere system. This is an abelian group, and the natural topology that it carries as a finite set or a Cantor set agrees with the subspace topology. Thus it is also compact. ∎
@@
Suppose $\Gamma$ is a rank $1$ graph. Let $f\in\mathrm{Diff}^{+}(M_{\Gamma})$ be so that $[f]\in\mathrm{ker}(\Psi)$. Then $f$ is isotopic to a product of sphere twists on the core spheres and the boundary components of the pieces of $M_{\Gamma}$. In particular, $\mathrm{ker}(\Psi)$ is a compact abelian subgroup of $\mathrm{Map}(M_{\Gamma})$.
@@
Our strategy will be to reduce down to the finite type case, which is covered by Proposition 3.3, and the rank 00 case, which is covered by Lemma 4.3. We may assume that $\Gamma$ is not of finite type, as otherwise we could apply Proposition 3.3 directly to show that $[f]\in\text{ker}(\Psi)$ is a product of sphere twists.
@@
By Proposition 3.3 and the discussion after it, we can isotope $f$ restricted to $K$ so that it is a composition of sphere twists on the core sphere of $K$ and sphere twists on spheres parallel to components of $\partial K$. By Lemmas 4.2 and 4.3, we can isotope $f$ on the components of $M_{\Gamma}\setminus K$ so that it a composition of sphere twists on the boundary components of pieces of $M_{\Gamma}$. The statement about $\mathrm{ker}(\Psi)$ follows. ∎
The previous proof really only relied on $\Gamma$ having finite rank, there was nothing special about rank $1$ in the argument.

### v2

@@ 1 Introduction
Essentially all of the work so far in studying mapping class groups of infinite type graphs has been in understanding the basic algebraic and topological properties of these groups, as well as their coarse geometry [algom-kfir_groups_2021] [domat_coarse_2023] [domat_generating_2023].
Given a locally finite connected graph $\Gamma$, we consider the associated doubled handlebody with punctures $M_{\Gamma}$. This is a $3$-manifold obtained by gluing two copies of a regular neighborhood of $\Gamma$ in ${\mathbb{R}}^{3}$ along their common surface boundary by the identity map (see Definition 2.3 for more details). Then the sphere complex $\mathcal{S}(M_{\Gamma})$ is the simplicial complex whose $k$-simplices are given by isotopy classes of $k+1$ essential non-peripheral smoothly embedded spheres in $M_{\Gamma}$ which can be realized disjointly. The mapping class group Map$(M_{\Gamma})$ of $M_{\Gamma}$, which is the group of orientation preserving diffeomorphisms of $M_{\Gamma}$ up to isotopy, acts naturally on this complex.
@@
Let $\Gamma$ be a locally finite connected graph, and let $M_{\Gamma}$ denote its associated doubled handlebody with punctures.
@@
    where $\mathrm{Twists}(M_{\Gamma})$ denotes the subgroup of $\mathrm{Map}(M_{\Gamma})$ generated by sphere twists on sphere systems of $M_{\Gamma}$. Further, as long as $\Gamma$ is not a sporadic graph, $\mathrm{Twists}(M_{\Gamma})$ is precisely the kernel of the action of $\mathrm{Map}(M_{\Gamma})$ on $\mathcal{S}(M_{\Gamma})$, so there is an induced faithful action of $\mathrm{Map}(\Gamma)$ on $\mathcal{S}(M_{\Gamma})$.
@@
Here, a sphere system is a smooth proper embedding of (possibly infinitely many) disjoint essential non-peripheral $2$-spheres in $M_{\Gamma}$. Unless otherwise stated, we assume that no pair of spheres are isotopic. A sphere in a $3$-manifold $M$ is essential if it does not bound a $3$-ball, and non-peripheral if it does not bound a once punctured $3$-ball and it is not isotopic to a boundary component. A sphere system is reduced if its complement is simply connected. A locally finite graph $\Gamma$ is sporadic if it is rank 00 with at most $4$ ends, a rank $1$ graph with at most $1$ end, or a compact rank $2$ graph. See the discussion before Conjecture 4.7.
@@
The group $\mathrm{Twists}(M_{\Gamma})$ is topologically isomorphic to $\Pi_{i=1}^{\text{rk}(\Gamma)}{\mathbb{Z}}/2$, with generating set given by sphere twists on a reduced sphere system (allowing for infinite products). In particular, $\mathrm{Twists}(M_{\Gamma})$ is compact and abelian.
@@
Finally, in Subsection 8.2 we consider a particular collection of graphs known as translatable graphs, analogous to the translatable surface of Schaffer-Cohen [schaffer-cohen_graphs_2024]. These can be thought of as graphs which can be formed by taking a collection of copies of a fixed graph arranged like ${\mathbb{Z}}$ with neighboring copies wedge summed along a point. See Definition 8.12 for a more precise definition. For a particular subclass of these graphs (those with tame end space, see Definition 8.16), we have the following.
@@
Acknowledgements: The author thanks his advisor Christopher J. Leininger for his support and guidance throughout this project. He’d also like to thank George Domat, Hannah Hoganson, and Mladen Bestvina for discussions about their work and for their support. Thanks to Thomas Hill, Michael Kopreski, Rebecca Rechkin, and George Shaji for helpful conversations on the topic while working on a project in parallel to this one. Finally, thanks to Sanghoon Kwak for helpful comments and corrections on an earlier version of the paper. In addition, the author acknowledges partial support from NSF grant DMS-1745670.
@@ 2.1 Space of ends
The collection of all such $U^{*}$ sets also forms a basis for a topology on $E(X)$.
@@ 2.2 Graphs and their doubled handlebodies
Let $\Gamma$ be a locally finite connected graph (every graph considered in this paper will be locally finite and connected, unless otherwise specified). Typically, $\Gamma$ will be noncompact. We consider the space of ends of $\Gamma$. The space of ends accumulated by loops of $\Gamma$, denoted $E_{\ell}(\Gamma)$, is the closed subset of $E(\Gamma)$ consisting of those ends whose neighborhoods in $\Gamma$ are all of infinite rank. Note that the elements of $E_{\ell}$ are called non-$\infty$-stable ends in [ayala_proper_1990] and ends accumulated by genus in [algom-kfir_groups_2021].
@@
The doubled handlebody (with punctures) $M_{\Gamma}$ of $\Gamma$ is the union of all the pieces of the vertices of $\Gamma$, with two boundary spheres quotiented together by an orientation reversing map if the corresponding vertices are connected by an edge in $\Gamma$.
@@
Typically, for consistency, when we need to take a compact exhaustion, we will assume that it is by unions of pieces.
@@
Fix a maximal sphere system $\mathcal{P}$ in $M_{\Gamma}$ (i.e. a sphere system so every essential non-peripheral sphere either intersects or is isotopic to some component of it). If $\Gamma$ is of infinite type, then $\mathcal{P}$ is infinite. Such a maximal system splits $M_{\Gamma}$ into a collection $\{P_{k}\}$ of copies of $M_{0,3}$ (sometimes one has to replace a boundary component of a copy of $M_{0,3}$ with a puncture corresponding to an end of $M_{\Gamma}$).
@@ 2.3 The Sphere Complex
An isotopy of a sphere in $M_{\Gamma}$ preserves the ends of $M_{\Gamma}$ it separates, so if two spheres are isotopic they must induce the same partition of $E(\Gamma)$. On the other hand, suppose $S$ and $S^{\prime}$ induce the same partition of $E(\Gamma)$. As $S$ and $S^{\prime}$ are compact, we can choose a finite number of disjoint spheres $T_{1},\ldots,T_{n}$ which are disjoint from $S$ and $S^{\prime}$ so that $S,S^{\prime}$ lie inside the submanifold bounded by $T_{1},\ldots,T_{n}$, and so that this submanifold is compact. As $S$ and $S^{\prime}$ induce the same partition of $E(\Gamma)$, they induce the same partition of the spheres $T_{1},\ldots,T_{n}$. But isotopy classes of spheres in manifolds homeomorphic to $M_{0,n}$ are determined by their partitions of boundary spheres, see Lemma 9 of [bering_iv_finite_2024]. ∎
Fix a sphere system $Q$ and an embedded sphere $S$ in normal form with respect to $Q$. We may identify an innermost disc $D$ of this intersection, which is a disc in a component of $Q$ so that $D$ does not contain any other component of $S\cap Q$. The surgery of $S$ along $D$ is the pair of spheres which result from cutting $S$ along $D$, and then attaching copies of $D$ to each of the two components. The resulting spheres are embedded and essential, though not always non-peripheral.
@@
Fix a basepoint $p\in\partial M_{n,1}$. We need one more auxiliary complex, which we will denote by $\widehat{\mathcal{S}}_{ns}(M_{n,1})$. It is the flag complex whose vertex set agrees with $\mathcal{S}_{ns}(M_{n,1})$. If $n\geq 3$, then there is an edge between two disjoint spheres $S$ and $S^{\prime}$ if $\pi_{1}(M_{n,1}\setminus(S\cup S^{\prime}),p)$ is nontrivial. If $n=2$, then there is an edge between disjoint spheres $S$ and $S^{\prime}$ if $\pi_{1}(M_{n,1}\setminus S,p)$ and $\pi_{1}(M_{n,1}\setminus S^{\prime},p)$ generate $\pi_{1}(M_{n,1},p)$. Note that for any $n\geq 2$, an edge in $\mathcal{S}_{ns}(M_{n,1})$ between vertices $S$ and $S^{\prime}$ is not in $\widehat{\mathcal{S}}_{ns}(M_{n,1})$ if and only if $S$, $S^{\prime}$, and $\partial M_{n,1}$ bound a copy of $M_{0,3}$.
@@
    Fix two disjoint vertices $S$ and $S^{\prime}$. If $S$ and $S^{\prime}$ are not joined by an edge in $\widehat{\mathcal{S}}_{ns}(M_{n,1})$, then it follows from the end of the paragraph after Lemma 2.15 that the complementary component of $S\cup S^{\prime}$ not containing $p$ is homeomorphic to $M_{n-1,2}$. As $n\geq 2$, there is a nonseparating sphere $S^{\prime\prime}$ in this component.
@@ 3.1 Mapping class groups of doubled handlebodies and graphs
where $\ell(t)\cdot x$ denotes the action of $\ell(t)\in\text{SO}(3)$ on $x\in S\cong S^{2}\subset{\mathbb{R}}^{3}$. We denote by $\text{Twists}(M_{\Gamma})$ the subgroup of $\text{Map}(M_{\Gamma})$ generated by compositions of the sphere twists on (potentially infinite) sphere systems of $M_{\Gamma}$. We equip $\text{Twists}(M_{\Gamma})$ with the subspace topology.
@@
We use the notation $\mathcal{V}_{K}$ for subgroups in both $\text{Map}(M_{\Gamma})$ and in $\text{Map}(\Gamma)$, but it should always be clear in context which one meant.
@@ 3.2 Constructing the homomorphism
We now construct the homomorphism from $\text{Map}(M_{\Gamma})$ to $\text{Map}(\Gamma)$ which appears in Theorem 1.1.
@@ 4 Short exact sequence
In this section, we give a proof of the first part of Theorem 1.1. The proof of the second part requires Theorem 1.2, and will be done in Section 5. We split the proof into two cases. The first case, where we assume $\Gamma$ has rank 00 or rank at least $2$, will be handled in Lemmas 4.2 and 4.3. The rank $1$ case will be handled in Lemma 4.4. For simplicity, let us assume that $\Gamma$ is in standard form. We can also assume that $\Gamma$ has no vertices of valence $1$, so in particular the boundary components of every piece of $M_{\Gamma}$ are essential (though potentially peripheral).
@@
We remark that there is a subtlety in Theorem 4.1, in that it deals with homotopies and isotopies of embeddings of spheres in $M$. This is as opposed to the definition of $\mathcal{S}(M_{\Gamma})$, which are isotopy classes of spheres considered as sets, instead of embeddings. As we will see, accounting for this difference is important in the proof of Theorem 1.1(1).
We will apply Theorem 4.1, along with the following lemma, to obtain Theorem 1.1 in the case when $\Gamma$ has rank 00 or rank at least $2$. This lemma is due to Laudenbach in the case that $\Gamma$ is compact [laudenbach_topologie_1974]. Our proof is based on the proof of the same result by Brendle–Broaddus–Putman. [brendle_mapping_2023].
@@
We lift to the universal cover $\widetilde{\Gamma}\subset\widetilde{M_{\Gamma}}$ with covering map $\pi:\widetilde{M_{\Gamma}}\to M_{\Gamma}$ and with a fixed lift $\widetilde{x_{0}}$ of $x_{0}$. Let $\tilde{f}:\widetilde{M_{\Gamma}}\to\widetilde{M_{\Gamma}}$ be the unique lift of $f$ fixing $\widetilde{x_{0}}$. We also denote by $\widetilde{i}:\widetilde{\Gamma}\to\widetilde{M_{\Gamma}}$ and $\widetilde{r}:\widetilde{M_{\Gamma}}\to\widetilde{\Gamma}$ the inclusion and retraction maps so that $\pi\circ\widetilde{i}=i\circ\pi$ and $\pi\circ\widetilde{r}=r\circ\pi$. To show that $f$ acts trivially on $\pi_{2}(M_{\Gamma},x_{0})$, it suffices to show that $\widetilde{f}$ acts trivially on $\pi_{2}(\widetilde{M_{\Gamma}},\widetilde{x_{0}})\cong H_{2}(\widetilde{M_{\Gamma}})$, by the Hurewicz Theorem.
@@
The next lemma gives a more explicit form of elements of $\text{ker}(\Psi)$. First, recall the core spheres of $M_{\Gamma}$, as defined in Subsection 2.4. This is a collection of nonseparating spheres in $M_{\Gamma}$, one for each piece of $M_{\Gamma}$ homeomorphic to $M_{1,s}$. This collection is not unique, but we will fix one from this point on.
@@
Let $\Gamma$ be a locally finite graph. Fix $f\in\text{Diff}^{+}(M_{\Gamma})$ acting trivially on $\pi_{2}(M,x_{0})$ for some $x_{0}\in M$. Then $f$ is isotopic to a product of sphere twists on the core spheres and the boundary components of the pieces of $M_{\Gamma}$. In particular, if $\Gamma$ has rank 00 or rank at least $2$, then $\mathrm{ker}(\Psi)$ consists of (possibly infinite) products of sphere twists on a spheres in a fixed sphere system.
@@
In this proof every sphere is actually implicitly an embedding $i:S^{2}\to M_{\Gamma}$. Thus when we say that two spheres are homotopic or isotopic, they are homotopic or isotopic as embeddings.
@@
Let $\mathcal{S}=\sqcup_{i}(\partial K_{i}\cup\mathcal{C}_{i})$. By Theorem 4.1, $f$ preserves the isotopy class of every component of $\mathcal{S}$. The goal is to produce an ambient isotopy of $M_{\Gamma}$ so that after applying this isotopy to $f$, every component of $\mathcal{S}$ is actually fixed by $f$. Note that we cannot apply Theorem 4.1 directly, as $\mathcal{S}$ has infinitely many elements. However, we will see that it can be inductively applied with care to achieve the desired result.
@@
Let $\widetilde{M_{\Gamma}}$ denote the universal cover of $M_{\Gamma}$ with $\pi:\widetilde{M_{\Gamma}}\to M_{\Gamma}$ the covering map. We can lift the homotopy between $S$ and $f_{n}(S)$ to one between lifts of these two spheres in $\widetilde{M_{\Gamma}}$. Let us denote these lifts by $T_{1}$ and $T_{2}$. By Theorem 4.1, they are isotopic as well. Thus these spheres lie in the same component $U$ of $\widetilde{M_{\Gamma}}\setminus\pi^{-1}(K_{n})$, or else there is no way they could induce the same partition of $E(\widetilde{M_{\Gamma}})$ (using the assumption that $K_{i}$ is not homeomorphic to $M_{0,2}$), contradicting Lemma 2.13. But then they also induce the same partition of $E(U)$, thinking of its boundary as an end. Again by Lemma 2.13, they are isotopic in $U$. This restricts to a homotopy from $T_{1}$ to $T_{2}$ whose image lies in $U$. Strictly speaking Lemma 2.13 only gives isotopy on the level of sets, instead of embeddings. However, the original homotopy between $T_{1}$ and $T_{2}$ is defined on embeddings, so it is clear that the homotopy between them whose image lies in $U$ can also be taken to be through embeddings.
@@
If we assume that $\Gamma$ has rank 00 or rank at least $2$, then by Lemma 4.2, every element of $\text{Ker}(\Psi)$ has a representative which acts trivially on $\pi_{2}(M_{\Gamma},x_{0})$. By the above, it follows then that $\text{Ker}(\Psi)$ is generated by compositions of sphere twists on the spheres of $\mathcal{S}$. ∎
@@
Suppose $\Gamma$ is a rank $1$ graph. Let $f\in\mathrm{Diff}^{+}(M_{\Gamma})$ be so that $[f]\in\mathrm{ker}(\Psi)$. Then $f$ is isotopic to a product of sphere twists on the core spheres and the boundary components of the pieces of $M_{\Gamma}$.
@@
Our strategy will be to reduce down to the finite type case, which is covered by Proposition 3.3, and the rank 00 case, which is covered by Lemma 4.3. As in the proof of Lemma 4.3, homotopies and isotopies between spheres are actually between embeddings of spheres. We may assume that $\Gamma$ is not of finite type, as otherwise we could apply Proposition 3.3 directly to show that $[f]\in\text{ker}(\Psi)$ is a product of sphere twists.
@@
By Proposition 3.3 and the discussion after it, we can isotope $f$ restricted to $K$ so that it is a composition of sphere twists on the core sphere of $K$ and sphere twists on spheres parallel to components of $\partial K$. By Lemmas 4.2 and 4.3, we can isotope $f$ on the components of $M_{\Gamma}\setminus K$ so that it a composition of sphere twists on the boundary components of pieces of $M_{\Gamma}$. ∎
The previous proof only relied on $\Gamma$ having finite rank, there was nothing special about rank $1$ in the argument.
We remark that, from now on, we will only consider spheres and sphere systems as embedded submanifolds, and forget about the actual embedding itself, unless stated otherwise. In particular, all homotopies and isotopies are for embedded submanifolds instead of for embeddings.
Recall that a locally finite graph $\Gamma$ is sporadic if it is rank 00 with at most $4$ ends, a rank $1$ graph with 00 or $1$ end(s), or a compact rank $2$ graph. These cases are explicitly excluded from the second part of Theorem 1.1(1). We will discuss after the proof of Theorem 1.1(1) why they need to be excluded.
Lemma 4.6.
Suppose $\Gamma$ is non-sporadic, and $[g]\in\text{Map}(M_{\Gamma})$ acts trivially on $\mathcal{S}(M_{\Gamma})$. Then it also fixes isotopy classes of embedded peripheral spheres of $M_{\Gamma}$.
Proof.
It suffices to show that $g$ fixes the ends of $M_{\Gamma}$, as a peripheral sphere is fixed if and only if the end it encloses is fixed.
If $\Gamma$ has rank 00, then by assumption $\Gamma$ has at least $5$ ends. Suppose $e\in E(M_{\Gamma})$ is such that $g(e)\neq e$. Let $S\in\mathcal{S}(M_{\Gamma})$ be a sphere containing $e$ and $g(e)$ on different sides, and so that the side containing $e$ contains at most $2$ ends. Then the side of $g(S)$ containing $g(e)$ contains at most $2$ ends, which is a contradiction by Lemma 2.13 as clearly $S$ and $g(S)$ do not induces the same partition on $E(\Gamma)$.
If the rank of $\Gamma$ is at least $1$, then we can assume it has at least $2$ ends (the rank $2$ case with $1$ end is trivial). Fix a sphere $S\in\mathcal{S}(M_{\Gamma})$ which cuts off a once punctured copy of $M_{1,1}$, where the puncture corresponds to an end $e$ of $M_{\Gamma}$. If $g(e)\neq e$, then it follows that $g(S)\neq S$ as $S$ and $g(S)$ partition the ends of $M_{\Gamma}$ differently which is impossible if they are isotopic. This is a contradiction, finishing the proof. ∎